problem:
(you have to use an equation to solve the problem)
the attendance at the ball game was 400 people. student tickets cost $2 and adult tickets cost $3. if $1050 was collected in ticket sales, how many of each type of ticket were sold?

Answer by LilSkittleMd(119)

The two equations are:
a+s=400
3a+2s=1050
Solve the first equation for a:
a=400-s
Substitute a into the second equation
3(400-s)+2s=1050
Distribute
1200-3s+2s=1050
1200-s=1050
-s=-150
s=150
Substitute the value of s into one of the original equations
a+150=400
a=250
There were 250 adult tickets sold and 150 student tickets sold